Coke and Pepsi compete in price competition. However, their products are differentiated – people can taste the difference!

Demand for Coke is given by Qc= 28-4Pc+0.1Pp.

Demand for Pepsi is given by Qp = 24-2Pp+0.1Pc.

The (constant) marginal cost for Coke is 3.

The (constant) marginal cost for Pepsi is 4.

What is the equilibrium price of Coke in this one-shot, simultaneous move game? Give your answer to two decimal places, for example, XX.XX.
Answer #1

1. Ans. Criven = - - - = 0 - 24. чPC + 0.1 Re_ Оe : 2ч - Pe F). Ра Marginal cost (coke) = 2 3 - Marginal cort (Pepsi = 4 Рң020+ 0.lle - PP - ④ Now Pulling value of 0 in ① we alti 20+ 0.1 [ 25.10.181 - PP 160 + 2.5+ o.oipe = 4PP 162-5+0.01pp = 32pp 1